Gans Incorporated developed a business strategy that uses stock options as a major compensation incentive for its top executives. On January 1, 2024, 20 million options were granted, each giving the executive owning them the right to acquire five $1 par common shares. The exercise price is the market price on the grant date—$10 per share. Options vest on January 1, 2028. They cannot be exercised before that date and will expire on December 31, 2030. The fair value of the 20 million options, estimated by an appropriate option pricing model, is $40 per option. Ignore income tax.

On March 1, 2028, when the market price of Gans Incorporated's stock was $14 per share, 3 million of the options were exercised. The journal entry to record this would include:
Answer #1

cash: 5 (five $1par common shares) *10 per share* 3million = 150 

pic: 3*40= 120

common stock 15(10+5)*$1 par= 15

150+120-15=255 

answer: a credit to paid in capital excess of par for $255 million

answered by: anonymous